# Context is Key: A Benchmark for Forecasting with Essential Textual Information

### v1

@@ B.1 Task: Constrained Predictions
Domain: Traffic
Context sources: Future information
Capabilities: Instruction Following Context: “Suppose that in the forecast, the values are bounded above by 11.88, the values are bounded below by 7.06.”
@@
Since we draw this series from the PeMS dataset (Chen et al., 2001), we tag its domain as “Traffic”. The context directly refers to the future, hence the context source is tagged as “Future information”. Finally, since the model is expected to obey the constraints in the context, we tag the evaluated capability as “Instruction following”.
@@ B.2 Task: Electrical Consumption Increase
Domain: Energy
Context sources: Future information, Covariate information
Capabilities: Instruction following, Retrieval from context
Context: “This is the electricity consumption recorded in Kilowatt (kW) in city A. A heatwave struck the city, which began on 2012-10-09 18:00:00 and lasted for approximately 3 hours, saw temperatures soar to unprecedented levels. According to the city’s electricity provider, power consumption during the peak of the heatwave reached approximately 5 times the typical usage for this time of year.”
@@
Since this series represents electricity consumption (Sec. 3.1), we tag it a coming from the “Energy” domain. The context sources for this task are twofold: the first context source is “Future information”, which represents knowledge of the five-fold increase in typical usage during the shock. The second source of context, “Covariate information”, represents the occurrence of a heatwave, which coincides with the timing and duration of the shock. The model must therefore interpret both the information on the magnitude of the shock from the future information, as well as the timing and duration of the sock from the covariate information. Together, these pieces of information enable an accurate forecast despite the lack of information about the shock in the task’s numerical history.
The skills for this task are tagged as “Instruction following” and “Retrieval from context”. While instruction following involves interpreting the context to include the shock in the prediction, the model must also retrieve from the context the relevant information, as there is unneeded information in the context as well: an accurate forecast does not require knowing that the temperature has reached unprecedented levels.
@@ B.3 Task: ATM Maintenance
Domain: Retail
Context sources: Intemporal information, Covariate information
Capabilities: Instruction following, Deductive reasoning
Context: “This is the number of cash withdrawals from an automated teller machine (ATM) in an arbitrary location in England. The ATM was under maintenance for 7 days, periodically every 14 days, starting from 1996-11-30 00:00:00. Assume that the ATM will not be in maintenance in the future.”
@@
Since this series represents ATM withdrawals, we tag it as “Retail”. The context includes information such as the location of the ATM, and therefore provides “Intemporal information”. As the maintenance frequency and duration is also described, the context sources include “Covariate information”.
This task is tagged with two capabilities. “Instruction following” is necessary because the model must assume that the ATM will not be in maintenance in the future. However, the model must use “Deductive reasoning” to determine what and when the impact of the maintenance was – reducing the number of withdrawals to 0 every 14 days –, and avoid including that pattern in the forecast. The RoI represents when the maintenance periods would have occurred in the forecast horizon, which is likely where forecasting models that do not leverage the context will forecast 0. While a quantitative forecasting model would find such a signal irresistible, context-aware models should avoid repeating the pattern in the forecast.
@@ B.4 Task: Montreal Fire High Season
Domain: Public Safety
Context sources: Intemporal information, Historical information
Capabilities: Deductive reasoning, Mathematical reasoning, Retrieval from memory
Context: “The Montreal Fire Department is in charge of responding to various kind of public safety incidents. This is the number of field fire incidents responded to by Montreal firefighters in the borough of Rivière-des-Prairies-Pointe-aux-Trembles. In other years, the yearly average number of incidents was 106 with the busiest month being June.”
@@
This task is tagged with many skills and involves several steps of interpretation to arrive at a reasonable forecast. We first note that the task requires “Retrieval from memory”: an important piece of information for this prediction is that winters in Montreal, a city in the northern hemisphere, are long and harsh, with temperatures reaching $-40^{\circ}$. Secondly, the task requires the model to use “Deductive reasoning” to deduce that, since temperatures are so cold during the winter months, fields are likely covered in snow and are rather unlikely to catch fire. Finally, the model can employ “Mathematical reasoning” to determine how many field fires are likely to occur on average in the forecast horizon, given the total number of field fires that have already blazed during the history.
Note that “Retrieval from memory” tasks do not explicitly ask the model to return information retrieved from memory; rather, we tag tasks as such because they cannot be solved without key information that is not present in the history or the context.
@@ B.5 Task: Solar Prediction
Domain: Climatology
Context sources: Intemporal information
Capabilities: Analogical reasoning, Deductive reasoning, Retrieval from memory
Context: “This series estimates the power production for a given day of a new solar power plant located in the state of Georgia, which has a climate similar to Alabama’s.”
@@
Without the “Intemporal information” that the context provides, it is quite possibly impossible to accurately forecast the parabola-like shape of the ground truth: the history contains very few defining characteristics, which makes it interchangeable with that of many potential processes and therefore many possible forecasts. The model must use “Deductive reasoning” to foresee this reversion to zero based on the fact that solar panels do not produce electricity at night.
However, the information in the context alone is not sufficient to provide an accurate forecast: nothing indicates the time at which production should peak. It must therefore rely on “Retrieval from memory” to retrieve information about Alabama’s climate and then “Analogical reasoning” to apply it to the present problem.
@@ B.6 Task: Speed From Load
Domain: Mechanics
Context sources: Causal information, Intemporal information, Covariate information
Capabilities: Causal reasoning, Mathematical reasoning, Instruction following
Context: “The wind tunnel is a chamber with one controllable fan that pushes air through it. We can control the load of the fan (corresponding to the duty cycle of the pulse-width-modulation signal) and measure its speed (in revolutions per minute). The fan is designed so its steady-state speed scales broadly linearly with the load. Unless completely powered off, the fan never operates below a certain speed, corresponding to a minimum effective load between 0.1 and 0.2. The task is to forecast the speed of the fan. The load is between 0 and 1. At full load (=1), the fan turns at a maximum speed of 3000 rpm. The load is set to: 0.0 until 05:47:09, 0.1 from 05:47:09 until 05:47:29, 0.0 from 05:47:29 until 05:48:01, 0.2 from 05:48:01 until 05:48:27, 0.1 from 05:48:27 until 05:48:49, 0.0 from 05:48:49 until 05:49:00.”
The “Speed From Load” task combines many different context sources and capabilities to produce a forecast of the revolutions per minute (RPM) of a fan in a wind tunnel based on its load. This task, based on the Causal Chambers dataset (Gamella et al., 2024), is therefore tagged as part of the “Mechanics” domain.
As the plot shows, producing an accurate forecast of the ground truth (orange line) from the numerical history alone (black line) is essentially impossible. However, the context of the task is quite rich: it provides “Intemporal information” on the nature of the task, such as the limits of the load and of the fan, “Covariate information” that describes the load during the history and future, as well as “Causal information” on the control that the load exerts on the fan, as well as the proportionality of their relationship.
To leverage the context requires multiple skills: firstly, “Instruction following” is necessary to understand that the task is to forecast the speed of the fan (as opposed to e.g. the load) and to apply the correct loads at the right moments. Secondly, the model must use “Causal reasoning” to understand that the changes in the load will directly impact the speed of the fan. Finally, the model must leverage “Mathematical reasoning” to calculate the speed of the fan as a function of the load.
@@ Appendix C Additional Results
C.1 Results partitioned by types of context
Table Tab. 2 provides a view of the results partitioned by the types of context. One can observe that Direct Prompt - Llama-3.1-405B-Instruct achieves the best performance at tasks where the context involves intemporal, future or covariate information, while GPT-4o has an upper hand at tasks involving historical context information. LLMP with Llama-3-70B-Instruct achieves the best performance in tasks that involve causal information in the context. This provides a view complementary to that of partitioning by model capabilities (as in Tab. 1), and emphasizes that no single model is the best at processing all types of context, leaving room for advancements in models in the future.
Table 2: Results on the CiK benchmark aggregated over all tasks and kinds of context. The first column shows the RCRPS averaged over all tasks. The second column shows the rank of each method w.r.t. other baselines, averaged over all tasks. The remaining columns show the average RCRPS stratified by context source. All averages are weighted according to the scheme described in Sec. 5.1 and accompanied by standard errors. Lower is better and the best means are in bold. * denotes models that do not use natural language context.
Model Average RCRPS Average Rank $\mathbf{c}_{I}$ $\mathbf{c}_{H}$ $\mathbf{c}_{F}$ $\mathbf{c}_{cov}$ $\mathbf{c}_{\text{causal}}$
Direct Prompt (ours)
     Llama-3.1-405B-Inst 0.159 $\pm$ 0.008 4.469 $\pm$ 0.192 0.174 $\pm$ 0.010 0.146 $\pm$ 0.001 0.085 $\pm$ 0.003 0.169 $\pm$ 0.010 0.398 $\pm$ 0.045
     Llama-3-70B-Inst 0.518 $\pm$ 0.030 10.406 $\pm$ 0.190 0.621 $\pm$ 0.042 0.308 $\pm$ 0.064 0.301 $\pm$ 0.033 0.452 $\pm$ 0.032 0.704 $\pm$ 0.056
     Llama-3-8B-Inst 1.647 $\pm$ 0.069 15.130 $\pm$ 0.171 2.355 $\pm$ 0.100 0.813 $\pm$ 0.115 1.332 $\pm$ 0.094 1.185 $\pm$ 0.087 2.041 $\pm$ 0.271
     Mixtral-8x7B-Inst 1.061 $\pm$ 0.058 13.381 $\pm$ 0.230 1.263 $\pm$ 0.082 0.561 $\pm$ 0.111 0.691 $\pm$ 0.094 0.724 $\pm$ 0.053 1.232 $\pm$ 0.121
     GPT-4o 0.276 $\pm$ 0.010 4.368 $\pm$ 0.149 0.220 $\pm$ 0.007 0.118 $\pm$ 0.001 0.108 $\pm$ 0.001 0.265 $\pm$ 0.012 0.858 $\pm$ 0.053
     GPT-4o-mini 0.353 $\pm$ 0.022 8.930 $\pm$ 0.177 0.474 $\pm$ 0.035 0.139 $\pm$ 0.002 0.141 $\pm$ 0.001 0.345 $\pm$ 0.030 0.644 $\pm$ 0.128
LLMP
     Llama-3-70B-Inst 0.550 $\pm$ 0.013 8.038 $\pm$ 0.205 0.455 $\pm$ 0.018 0.516 $\pm$ 0.028 0.690 $\pm$ 0.018 0.588 $\pm$ 0.018 0.392 $\pm$ 0.028
     Llama-3-70B 0.237 $\pm$ 0.006 6.560 $\pm$ 0.254 0.213 $\pm$ 0.005 0.121 $\pm$ 0.008 0.233 $\pm$ 0.012 0.198 $\pm$ 0.004 0.360 $\pm$ 0.011
     Llama-3-8B-Inst 0.484 $\pm$ 0.010 9.457 $\pm$ 0.166 0.477 $\pm$ 0.013 0.161 $\pm$ 0.006 0.264 $\pm$ 0.003 0.316 $\pm$ 0.008 0.878 $\pm$ 0.035
     Llama-3-8B 0.313 $\pm$ 0.023 9.499 $\pm$ 0.323 0.334 $\pm$ 0.035 0.123 $\pm$ 0.004 0.232 $\pm$ 0.012 0.291 $\pm$ 0.031 0.739 $\pm$ 0.134
     Mixtral-8x7B-Inst 0.264 $\pm$ 0.004 8.496 $\pm$ 0.256 0.242 $\pm$ 0.007 0.173 $\pm$ 0.004 0.268 $\pm$ 0.009 0.220 $\pm$ 0.002 0.437 $\pm$ 0.007
     Mixtral-8x7B 0.262 $\pm$ 0.008 8.619 $\pm$ 0.208 0.250 $\pm$ 0.008 0.119 $\pm$ 0.003 0.254 $\pm$ 0.013 0.229 $\pm$ 0.007 0.457 $\pm$ 0.011
Multimodal Models
     UniTime 0.371 $\pm$ 0.002 13.495 $\pm$ 0.091 0.455 $\pm$ 0.002 0.154 $\pm$ 0.000 0.226 $\pm$ 0.003 0.396 $\pm$ 0.001 0.422 $\pm$ 0.001
     TimeLLM 0.476 $\pm$ 0.001 16.662 $\pm$ 0.075 0.517 $\pm$ 0.002 0.183 $\pm$ 0.000 0.376 $\pm$ 0.002 0.446 $\pm$ 0.001 0.482 $\pm$ 0.001
TS Foundation Models*
     Lag-Llama 0.329 $\pm$ 0.004 13.157 $\pm$ 0.235 0.333 $\pm$ 0.005 0.167 $\pm$ 0.005 0.277 $\pm$ 0.006 0.301 $\pm$ 0.004 0.495 $\pm$ 0.014
     Chronos 0.326 $\pm$ 0.002 11.962 $\pm$ 0.139 0.314 $\pm$ 0.002 0.179 $\pm$ 0.003 0.316 $\pm$ 0.002 0.252 $\pm$ 0.002 0.460 $\pm$ 0.004
     TimeGEN 0.354 $\pm$ 0.000 14.345 $\pm$ 0.090 0.333 $\pm$ 0.000 0.177 $\pm$ 0.000 0.348 $\pm$ 0.000 0.291 $\pm$ 0.000 0.474 $\pm$ 0.000
     Moirai 0.520 $\pm$ 0.006 12.458 $\pm$ 0.266 0.596 $\pm$ 0.009 0.140 $\pm$ 0.001 0.364 $\pm$ 0.002 0.510 $\pm$ 0.008 0.438 $\pm$ 0.011
Statistical Models*
     ARIMA 0.480 $\pm$ 0.006 12.320 $\pm$ 0.180 0.565 $\pm$ 0.010 0.200 $\pm$ 0.007 0.307 $\pm$ 0.003 0.390 $\pm$ 0.006 0.440 $\pm$ 0.011
     ETS 0.522 $\pm$ 0.009 14.310 $\pm$ 0.203 0.627 $\pm$ 0.014 0.362 $\pm$ 0.014 0.323 $\pm$ 0.008 0.401 $\pm$ 0.010 0.508 $\pm$ 0.017
     Exp-Smoothing 0.603 $\pm$ 0.013 14.936 $\pm$ 0.132 0.700 $\pm$ 0.020 0.493 $\pm$ 0.016 0.438 $\pm$ 0.009 0.492 $\pm$ 0.017 0.827 $\pm$ 0.060
C.2 Extended results on all models
Table 3: Extended results on the CiK benchmark aggregated over all tasks. The first column shows the RCRPS averaged over all tasks. The second column shows the rank of each method w.r.t. other baselines, averaged over all tasks. All averages are weighted according to the scheme described in Sec. 5.1 and accompanied by standard errors. Lower is better and the best means are in bold.

### v2

@@ B.1 Task: Constrained Predictions
Domain: Traffic Context types: Future information Context: “Suppose that in the forecast, the values are bounded above by 11.88, the values are bounded below by 7.06.” [Uncaptioned image]
@@
Since we draw this series from the PeMS dataset (Chen et al., 2001), we tag its domain as “Traffic”. The context directly refers to the future, hence the context source is tagged as “Future information”.
@@ B.2 Task: Electrical Consumption Increase
Domain: Energy Context types: Future information, Covariate information Context: “This is the electricity consumption recorded in Kilowatt (kW) in city A. A heatwave struck the city, which began on 2012-10-09 18:00:00 and lasted for approximately 3 hours, saw temperatures soar to unprecedented levels. According to the city’s electricity provider, power consumption during the peak of the heatwave reached approximately 5 times the typical usage for this time of year.” [Uncaptioned image]
@@
Since this series represents electricity consumption (Sec. 3.1), we tag it a coming from the “Energy” domain. The context types for this task are twofold: the first context source is “Future information”, which represents knowledge of the five-fold increase in typical usage during the shock. The second source of context, “Covariate information”, represents the occurrence of a heatwave, which coincides with the timing and duration of the shock. The model must therefore interpret both the information on the magnitude of the shock from the future information, as well as the timing and duration of the sock from the covariate information. Together, these pieces of information enable an accurate forecast despite the lack of information about the shock in the task’s numerical history.
@@ B.3 Task: ATM Maintenance
Domain: Retail Context types: Intemporal information, Covariate information Context: “This is the number of cash withdrawals from an automated teller machine (ATM) in an arbitrary location in England. The ATM was under maintenance for 7 days, periodically every 14 days, starting from 1996-11-30 00:00:00. Assume that the ATM will not be in maintenance in the future.” [Uncaptioned image]
@@
Since this series represents ATM withdrawals, we tag it as “Retail”. The context includes information such as the location of the ATM, and therefore provides “Intemporal information”. As the maintenance frequency and duration is also described, the context types include “Covariate information”.
The RoI represents when the maintenance periods would have occurred in the forecast horizon, which is likely where forecasting models that do not leverage the context will forecast 0. While a quantitative forecasting model would find such a signal irresistible, context-aware models should avoid repeating the pattern in the forecast.
@@ B.4 Task: Montreal Fire High Season
Domain: Public Safety Context types: Intemporal information, Historical information Context: “The Montreal Fire Department is in charge of responding to various kind of public safety incidents. This is the number of field fire incidents responded to by Montreal firefighters in the borough of Rivière-des-Prairies-Pointe-aux-Trembles. In other years, the yearly average number of incidents was 106 with the busiest month being June.” [Uncaptioned image]
@@ B.5 Task: Solar Prediction
Domain: Climatology Context types: Intemporal information Context: “This series estimates the power production for a given day of a new solar power plant located in the state of Georgia, which has a climate similar to Alabama’s.” [Uncaptioned image]
@@
Without the “Intemporal information” that the context provides, accurate forecasts of the parabola-like shape of the ground truth are unlikely: the history contains very few defining characteristics, which makes it interchangeable with that of many potential processes and therefore many possible forecasts.
@@ B.6 Task: Speed From Load
Domain: Mechanics Context types: Causal information, Intemporal information, Covariate information Context: “The wind tunnel is a chamber with one controllable fan that pushes air through it. We can control the load of the fan (corresponding to the duty cycle of the pulse-width-modulation signal) and measure its speed (in revolutions per minute). The fan is designed so its steady-state speed scales broadly linearly with the load. Unless completely powered off, the fan never operates below a certain speed, corresponding to a minimum effective load between 0.1 and 0.2. The task is to forecast the speed of the fan. The load is between 0 and 1. At full load (=1), the fan turns at a maximum speed of 3000 rpm. The load is set to: 0.0 until 05:47:09, 0.1 from 05:47:09 until 05:47:29, 0.0 from 05:47:29 until 05:48:01, 0.2 from 05:48:01 until 05:48:27, 0.1 from 05:48:27 until 05:48:49, 0.0 from 05:48:49 until 05:49:00.” [Uncaptioned image]
The “Speed From Load” task combines many different context types and capabilities to produce a forecast of the revolutions per minute (RPM) of a fan in a wind tunnel based on its load. This task, based on the Causal Chambers dataset (Gamella et al., 2024), is therefore tagged as part of the “Mechanics” domain.
As the plot shows, producing an accurate forecast of the ground truth (orange line) from the numerical history alone (black line) is essentially impossible. However, the context of the task is quite rich: it provides “Intemporal information” on the nature of the task, such as the limits of the load and of the fan, “Covariate information” that describes the load during the history and future, as well as “Causal information” on the control that the load exerts on the fan, as well as the proportionality of their relationship. With this information, it is possible to forecast the series nearly perfectly, excepting some noise.
@@ Appendix C Additional Results
The full unaggregated results from our benchmark can be found at the following location: https://anonymous.4open.science/r/context-is-key-forecasting-E391/results/results_complete.csv.
C.1 Extended results on all models
Table 2: Results of all models in the CiK benchmark aggregated over all tasks. The first column shows the RCRPS averaged over all tasks. The second column shows the rank of each method w.r.t. other baselines, averaged over all tasks. All averages are weighted according to the scheme described in Sec. 5.1 and accompanied by standard errors. Lower is better and the best means are in bold.
@@ C.1 Extended results on all models
direct prompt
Llama-3.1-405B-Inst
0.159 $\pm$ 0.008 7.453 $\pm$ 0.587
Llama-3-70B-Inst
0.237 $\pm$ 0.006 11.966 $\pm$ 0.675
Llama-3-8B-Inst
0.313 $\pm$ 0.023 18.354 $\pm$ 0.957
Mixtral-8x7B-Inst
0.523 $\pm$ 0.023 31.392 $\pm$ 0.391
Qwen-2.5-7B-Inst
0.289 $\pm$ 0.003 21.785 $\pm$ 0.725
Qwen-2.5-1.5B-Inst
0.575 $\pm$ 0.014 30.636 $\pm$ 0.839
Qwen-2.5-0.5B-Inst
0.464 $\pm$ 0.012 26.255 $\pm$ 0.517
GPT-4o
0.276 $\pm$ 0.010 8.607 $\pm$ 0.408
GPT-4o-mini
0.353 $\pm$ 0.022 16.469 $\pm$ 0.470
llmp
Llama-3-70B-Inst
0.550 $\pm$ 0.013 17.043 $\pm$ 0.539
Llama-3-70B
0.237 $\pm$ 0.006 11.966 $\pm$ 0.675
Llama-3-8B-Inst
0.484 $\pm$ 0.010 18.221 $\pm$ 0.444
Llama-3-8B
0.313 $\pm$ 0.023 18.354 $\pm$ 0.957
Mixtral-8x7B-Inst
0.264 $\pm$ 0.004 15.213 $\pm$ 0.609
Mixtral-8x7B
0.262 $\pm$ 0.008 15.903 $\pm$ 0.483
Qwen-2.5-7B-Inst
1.980 $\pm$ 0.027 42.499 $\pm$ 0.705
Qwen-2.5-7B
0.904 $\pm$ 0.037 35.484 $\pm$ 0.969
Qwen-2.5-1.5B-Inst
2.153 $\pm$ 0.027 47.551 $\pm$ 0.845
Qwen-2.5-1.5B
1.731 $\pm$ 0.036 42.323 $\pm$ 0.520
Qwen-2.5-0.5B-Inst
1.938 $\pm$ 0.024 47.515 $\pm$ 0.582
Qwen-2.5-0.5B
1.991 $\pm$ 0.024 45.717 $\pm$ 0.806
Multimodal Models
UniTime
0.370 $\pm$ 0.001 33.379 $\pm$ 0.150
Time-LLM (ETTh1)
0.476 $\pm$ 0.001 41.640 $\pm$ 0.143
Without Context
direct prompt
Llama-3.1-405B-Inst
0.473 $\pm$ 0.005 32.452 $\pm$ 0.282
Llama-3-70B-Inst
0.470 $\pm$ 0.008 33.490 $\pm$ 0.348
Llama-3-8B-Inst
0.601 $\pm$ 0.006 39.660 $\pm$ 0.385
Mixtral-8x7B-Inst
0.712 $\pm$ 0.021 44.241 $\pm$ 0.354
Qwen-2.5-7B-Inst
0.420 $\pm$ 0.022 35.119 $\pm$ 0.509
Qwen-2.5-1.5B-Inst
0.450 $\pm$ 0.007 34.473 $\pm$ 0.589
Qwen-2.5-0.5B-Inst
0.305 $\pm$ 0.007 25.475 $\pm$ 0.328
GPT-4o
0.441 $\pm$ 0.008 29.853 $\pm$ 0.364
GPT-4o-mini
0.423 $\pm$ 0.006 34.036 $\pm$ 0.265
llmp
Llama-3-70B-Inst
0.378 $\pm$ 0.004 24.662 $\pm$ 0.446
Llama-3-70B
0.312 $\pm$ 0.006 20.971 $\pm$ 0.433
Llama-3-8B-Inst
0.503 $\pm$ 0.009 29.624 $\pm$ 0.415
Llama-3-8B
0.345 $\pm$ 0.003 23.987 $\pm$ 0.359
Mixtral-8x7B-Inst
0.383 $\pm$ 0.015 23.530 $\pm$ 0.436
Mixtral-8x7B
0.306 $\pm$ 0.007 21.667 $\pm$ 0.471
Qwen-2.5-7B-Inst
1.023 $\pm$ 0.026 37.605 $\pm$ 1.169
Qwen-2.5-7B
0.732 $\pm$ 0.030 35.795 $\pm$ 1.092
Qwen-2.5-1.5B-Inst
1.515 $\pm$ 0.033 44.193 $\pm$ 1.030
Qwen-2.5-1.5B
1.069 $\pm$ 0.028 38.519 $\pm$ 1.037
Qwen-2.5-0.5B-Inst
1.318 $\pm$ 0.037 41.706 $\pm$ 0.734
Qwen-2.5-0.5B
1.819 $\pm$ 0.027 45.104 $\pm$ 0.690
Multimodal Models
UniTime
0.405 $\pm$ 0.002 35.049 $\pm$ 0.175
Time-LLM (ETTh1)
0.458 $\pm$ 0.002 40.647 $\pm$ 0.155
TS Foundation Models
Lag-Llama
0.329 $\pm$ 0.004 29.089 $\pm$ 0.782
Chronos-Tiny
0.328 $\pm$ 0.001 26.276 $\pm$ 0.416
Chronos-Mini
0.341 $\pm$ 0.001 27.559 $\pm$ 0.397
Chronos-Small
0.328 $\pm$ 0.002 25.213 $\pm$ 0.350
Chronos-Base
0.672 $\pm$ 0.003 29.216 $\pm$ 0.352
Chronos-Large
0.326 $\pm$ 0.002 24.463 $\pm$ 0.381
TimeGEN
0.354 $\pm$ 0.000 34.433 $\pm$ 0.163
Moirai-Small
0.565 $\pm$ 0.031 33.879 $\pm$ 0.419
Moirai-Base
0.624 $\pm$ 0.013 33.144 $\pm$ 0.380
Moirai-Large
0.520 $\pm$ 0.006 27.013 $\pm$ 0.853
Statistical Models
ARIMA
0.480 $\pm$ 0.006 25.571 $\pm$ 0.440
ETS
0.522 $\pm$ 0.009 31.842 $\pm$ 0.592
Exp-Smoothing
0.603 $\pm$ 0.013 34.049 $\pm$ 0.335
Tab. 2 provides the extended results with all models evaluated on the CiK benchmark, aggregated over all tasks.
C.2 Full results partitioned by types of context
Table 3: Results of all models on the CiK benchmark. Starting from the left, the first column shows the RCRPS averaged over all tasks. The second column shows the rank of each method w.r.t. other baselines, averaged over all tasks. The remaining columns show the average RCRPS stratified by types of context. All averages are weighted according to the scheme described in Sec. 5.1 and accompanied by standard errors. Lower is better and the best averages are in bold.
{tabu}
lccccccc Model Average RCRPS Average Rank Intemporal Information Historical Information Future Information Covariate Information Causal Information
With Context